# Non-Equilibrium Steady States and Quantum Chaos in a three-site Driven-Dissipative Bose-Hubbard Chains base on Self-Consistent Mean-Field Approach

### v5

@@ Abstract
We investigate the non-equilibrium dynamics and steady-state properties of a driven-dissipative Bose-Hubbard chain using a self-consistent Gutzwiller mean-field (GMF) approach. By employing a robust Picard iteration scheme, we solve the non-linear master equation for the non-equilibrium steady state (NESS) in the presence of strong Kerr nonlinearity. We identify two distinct dynamical regimes governed by the interplay between coherent drive, dissipation, and interaction: a regular quasilinear regime and a chaotic regime. Linear stability analysis reveals that the transition to the chaotic regime is triggered by parametric instabilities arising from the drive-induced coherence. Furthermore, we characterize the onset of quantum chaos by calculating the out-of-time-order correlator (OTOC). Our results show that in the strong coupling regime, the OTOC exhibits rapid exponential growth and saturation, providing a clear signature of information scrambling in this open quantum system. The proposed numerical framework offers an efficient pathway to explore many-body correlations in larger photonic lattices.
@@ IV Numerical Solution for NESS in the Non-Linear Regime ($U\neq 0$)
Using Roth’s Lemma ($\text{vec}(ABC)=(C^{T}\otimes A)\text{vec}(B)$), the Liouvillian reads
@@ V OTOC
In the Gutzwiller mean-field framework linearized around the NESS, this OTOC is related to the squared magnitude of the retarded Green’s function (propagator) governing the fluctuations $\delta\hat{a}$:
$D_{1,L}(\tau)\approx|G^{R}_{L,1}(\tau)|^{2}=\left|\left(e^{\mathbf{L}_{\text{chain}}\tau}\right)_{L,1}\right|^{2},$ (37)
where $\mathbf{L}_{\text{chain}}$ is the collective effective Liouvillian for the coupled chain fluctuations. Exponential growth of $D_{1,L}(\tau)$ indicates dynamical instability and the onset of chaos.
Next we consider the out-of-time-ordered correlator (OTOC)
$\mathbf{D}_{ij}(\tau)=\text{Tr}\left[\hat{\rho}_{SS}[a_{i}(\tau),a^{{\dagger}}_{j}(0)]^{2}\right]=\langle[\hat{a}_{i}(\tau),\hat{a}_{j}^{\dagger}(0)]^{\dagger}[\hat{a}_{i}(\tau),\hat{a}_{j}^{\dagger}(0)]\rangle_{SS}$ (38)
$=\langle\hat{a}_{j}(0)\hat{a}_{i}^{{\dagger}}(\tau)\hat{a}_{i}(\tau)\hat{a}_{j}^{\dagger}(0)-\hat{a}_{j}(0)\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{j}^{\dagger}(0)\hat{a}_{i}(\tau)-\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{j}(0)\hat{a}_{i}(\tau)\hat{a}_{j}^{\dagger}(0)+\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{j}(0)\hat{a}_{j}^{\dagger}(0)\hat{a}_{i}(\tau)\rangle$
$=\frac{2\langle\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{i}(\tau)\rangle\langle\hat{a}_{j}^{\dagger}(0)\hat{a}_{j}(0)\rangle-2\text{Re}(\langle\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{j}(0)\hat{a}_{i}(\tau)\hat{a}_{j}^{\dagger}(0)\rangle)}{\langle\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{i}(\tau)\rangle\langle\hat{a}_{j}^{\dagger}(0)\hat{a}_{j}(0)\rangle}$
$=2\left(1-\text{Re}\left[\frac{\langle a_{i}^{\dagger}(\tau)a_{j}a_{i}(\tau)a^{{\dagger}}_{j}\rangle_{SS}}{\langle a_{i}^{\dagger}a_{i}\rangle_{SS}\langle a_{j}a^{{\dagger}}_{j}\rangle_{SS}}\right]\right)$
@@
In the weak drive/linear limit, commutators decay or remain constant. $D_{1,L}(\tau\to\infty)\approx 0$. While in the strong coupling regime ($U\gg\gamma$), the non-linearity scrambles local information globally. The OTOC exhibits characteristic exponential growth at early times, $D(\tau)\sim e^{\lambda_{L}\tau}$ (where $\lambda_{L}$ is the quantum Lyapunov exponent), before saturating to a finite value $O(1)$ determined by the local Hilbert space dimension. This saturation indicates that information about the drive at site 1 has been effectively "forgotten" or scrambled across the entire many-body Hilbert space of the chain.
@@
Fig.2 show the OTOC as a function of time in unit of inverse hopping amplitude ($1/J$; such that the value $t=1.0$ corresponds to the characteristic time required for a particle to tunnel between neighboring sites). for site 3 (drain site) is of the strong dissipative and quasilinear regime, Although quantum information propagates to this site as indicated by the initial rise, it is quickly lost to the environment due to strong coupling with the bath. The vanishing saturation value is consistent with a strong dissipation or quasilinear regime, where information is not scrambled locally but rather "flushed out" before it can accumulate.

### v6

@@ Abstract
We investigate the non-equilibrium dynamics and steady-state properties of a driven-dissipative Bose-Hubbard chain using a self-consistent Gutzwiller mean-field (GMF) approach. By employing a robust Picard iteration scheme, we solve the non-linear master equation for the non-equilibrium steady state (NESS) in the presence of strong Kerr nonlinearity. We identify two distinct dynamical regimes governed by the interplay between coherent drive, dissipation, and interaction: a regular quasilinear regime and a chaotic regime. Linear stability analysis reveals that the transition to the chaotic regime is triggered by parametric instabilities arising from the drive-induced coherence. Importantly, we systematically study the intermediate stable nonlinear regime between the linear regime and chaotic regime, whose most significant character is the restored U(1) symmetry and the operatator OTOC whihc decays to zero at long time. Furthermore, we characterize the onset of quantum chaos by calculating the out-of-time-order correlator (OTOC). Our results show that in the strong coupling regime, the OTOC exhibits rapid exponential growth and saturation, providing a clear signature of information scrambling in this open quantum system. The proposed numerical framework offers an efficient pathway to explore many-body correlations in larger photonic lattices.
@@ IV Numerical Solution for NESS in the Non-Linear Regime ($U\neq 0$)
Using column-stacking Roth’s Lemma ($\text{vec}(ABC)=(C^{T}\otimes A)\text{vec}(B)$), the Liouvillian reads
@@ V OTOC
V.1 local OTOC
In the Gutzwiller mean-field framework linearized around the NESS, this OTOC is related to the squared magnitude of the retarded Green’s function (propagator) governing the fluctuations $\delta\hat{a}$
$D_{1,L}(t)=|G^{R}_{L,1}(t)|^{2}=\left|\left(e^{\mathbf{L}_{\text{tot}}t}\right)_{L,1}\right|^{2},$ (37)
where the retarded Green function reads $G^{R}_{L,1}(t)=\text{Tr}\{\hat{a}_{L}e^{\mathcal{L}t}[\hat{a}_{1}^{\dagger},\rho_{SS}]\}$. To find the change in $\langle\hat{a}_{L}\rangle$ due to a perturbation at site 1, we consider the vectorized Master Equation $\frac{d}{dt}|\rho\rangle\rangle=\mathbf{L}_{tot}|\rho\rangle\rangle$. A perturbation at site 1 at $t=0$ creates a deviation $\delta\rho(0)=[\hat{a}_{1}^{\dagger},\hat{\rho}_{SS}]$ such that the instantaneous state reads $\hat{\rho}(0^{+})=\hat{\rho}_{SS}+\epsilon\delta\hat{\rho}(0)$. Then the linear response reads
$\delta\langle\hat{a}_{L}(t)\rangle=\text{Tr}\{\hat{a}_{L}e^{\mathcal{L}t}\delta\hat{\rho}(0)\}=\langle\langle a_{L}|e^{\mathbf{L}_{tot}t}\delta|\rho(0)\rangle\rangle,$ (38)
The sensitivity is the projection of the evolved deviation onto the target operator
$D_{1,L}(t)=\left|\langle\langle a_{L}|e^{\mathbf{L}_{tot}t}\delta|\rho(0)\rangle\rangle\right|^{2}=\left|\text{Tr}\{\hat{a}_{L}e^{\mathcal{L}t}[\hat{a}_{1}^{\dagger},\hat{\rho}_{SS}]\}\right|^{2}$ (39)
$\mathbf{L}_{\text{tot}}$ is the collective effective Liouvillian ($(N^{N_{max}})^{2}\times(N^{N_{max}})^{2}$) for the full system reads
$\mathbf{L}_{\text{tot}}=-i(\mathbb{I}\otimes\hat{H}_{tot}-\hat{H}_{tot}^{T}\otimes\mathbb{I})+\sum_{l}\gamma_{l}\left(\hat{a}_{l}^{*}\otimes\hat{a}_{l}-\frac{1}{2}\mathbf{I}\otimes\hat{a}_{l}^{\dagger}\hat{a}_{l}-\frac{1}{2}\hat{a}_{l}^{T}\hat{a}_{l}^{*}\otimes\mathbf{I}\right).$ (40)
where $\hat{a}_{l}(0)=\hat{I}_{1}\otimes\hat{I}_{2}\otimes\dots\otimes\hat{a}_{l}\otimes\dots\otimes\hat{I}_{L}$ with $\hat{I}$ the $4\times 4$ identity matrix. In the expression $|\hat{a}_{1}(t)\rangle\rangle=e^{\mathbf{L}_{Heis}t}|\hat{a}_{1}(0)\rangle\rangle$, the matrix exponential acts as a rotation operator in the 4096-dimensional basis space. For example, the non-zero off-diagonal blocks generated by $J$ allow the weight of the operator $\hat{a}_{1}$ to leak into the indices corresponding to $\hat{a}_{2}$, $\hat{a}_{3}$, $\hat{n}_{1}\hat{a}_{2}$, etc., simulating the spatial spread of information. Eventually, the evolved operator $\hat{a}_{i}(t)$ has non-zero components across almost all $D^{2}$ basis operators. Here in the Lindblad framework, the adjoint Liouvillian $\mathcal{L}^{\dagger}$
$\frac{d}{dt}\hat{a}_{l}(t)=\mathcal{L}^{\dagger}\hat{a}_{l}(t)\implies\frac{d}{dt}|\hat{a}_{l}(t)\rangle\rangle=\mathbf{L}_{Heis}|\hat{a}_{l}(t)\rangle\rangle$ (41)
$L_{Heis}$ accounts for the global system ($(N^{L})^{2}\times(N^{L})^{2}=4096^{2}$). It contains the hopping $J$ as internal coupling elements, allowing site 1 information to leak into site 3. While the local one $\mathbf{L}_{local}$ ($N^{2}\times N^{2}=16^{2}$ matrix) replaces hopping with a static drive, severing the dynamic causal link between sites.
@@
$\frac{d}{dt}|\hat{a}_{l}(t)\rangle\rangle=\mathbf{L}_{local}^{\dagger}|\hat{a}_{l}(t)\rangle\rangle,$ (42)
$|\hat{a}_{l}(t)\rangle\rangle=e^{\mathbf{L}_{local}^{\dagger}\tau}|\hat{a}_{l}(0)\rangle\rangle$
where $\mathbf{L}_{local}^{{\dagger}}$ is the local Heisenberg-Liouvillian, $\mathbf{L}_{local}$ is the local Schrodinger Liouvillian ($N^{2}\times N^{2}$) is constructed using the Kronecker product of local operators $\hat{H}_{l}^{\text{eff}}(\Psi)$ and $\hat{a}$
$\mathbf{L}_{local}=-i(\mathbb{I}\otimes\hat{H}_{l}^{\text{eff}}(\Psi)-(\hat{H}_{l}^{\text{eff}}(\Psi))^{T}\otimes\mathbf{I})+\gamma\left(\hat{a}_{l}^{*}\otimes\hat{a}_{l}-\frac{1}{2}\mathbf{I}\otimes\hat{a}_{l}^{\dagger}\hat{a}_{l}-\frac{1}{2}\hat{a}_{l}^{T}\hat{a}_{l}^{*}\otimes\mathbf{I}\right)$ (43)
where we use the conventions $(A\otimes B)\text{vec}(X)=\text{vec}(AXB^{T})$, $\text{vec}(AXB)=(A\otimes B^{T})\text{vec}(X)$ $L^{*}\rho L^{T}=(L\rho L^{\dagger})^{T}$, $\rho H^{T}-H^{T}\rho=-(H\rho-\rho H)^{T}$. In GMF approximation, $\hat{H}_{l}^{\text{eff}}(\Psi)$ (Eq.(28)) contains the static fields $\Psi_{l\pm 1}$ from neighbors where $\Psi_{l\pm 1}=\langle\hat{a}_{l\pm 1}\rangle_{SS}$ are fixed complex numbers obtained from Picard iteration, effectively treating the hopping as a constant local drive. The OTOC at site $l$ is calculated using the commutator evaluated in the local Hilbert space
$D_{l}(t)=\text{Tr}\left\{\hat{\rho}_{SS,l}\left[\hat{a}_{l}(t),\hat{a}_{l}^{\dagger}(0)\right]^{\dagger}\left[\hat{a}_{l}(t),\hat{a}_{l}^{\dagger}(0)\right]\right\}$ (44)
where $\hat{a}_{l}(t)$ is reconstructed from the above vectorized evolution. This single-site OTOC represents the local decoherence and sensitivity to local perturbations under a static mean-field environment. At $t=0$, the cross term $\hat{a}_{i}(0)\hat{a}_{i}^{\dagger}(t)\hat{a}_{i}^{\dagger}(0)\hat{a}_{i}(t)-\hat{a}_{i}^{\dagger}(t)\hat{a}_{i}(0)\hat{a}_{i}(t)\hat{a}_{i}^{\dagger}(0)=-2\langle\hat{n}_{i}^{2}+\hat{n}_{i}\rangle_{SS}$ is at its maximal magnitude with perfect phase correlation. For $t>0$, chaotic dynamics smear the phase. The cross term decays from its initial value toward zero. The incoherent terms at $t=0$ reads $\langle\hat{a}_{i}(0)\hat{a}_{i}^{{\dagger}}(t)\hat{a}_{i}(t)\hat{a}_{i}^{\dagger}(0)+\hat{a}_{i}^{\dagger}(t)\hat{a}_{i}(0)\hat{a}_{i}^{\dagger}(0)\hat{a}_{i}(t)\rangle=\langle 2\hat{n}^{2}+2\hat{n}+1\rangle_{SS}$. Thus in an infinite-dimensional Hilbert space, the commutation relation $\hat{a},\hat{a}^{\dagger}]=\hat{I}$ holds exactly and the single site OTOC reads $\mathbf{D}_{ii}(0)=\text{Tr}\{\hat{\rho}_{SS;i}[\hat{a}_{i},\hat{a}_{i}^{\dagger}]^{\dagger}[\hat{a}_{i},\hat{a}_{i}^{\dagger}]\}=\text{Tr}\{\hat{\rho}_{SS;i}\cdot\hat{I}\}=1$. While for simulation shown in Fig.2, we consider a truncated Fock basis of dimension $N=N_{max}+1=4$, thus $\hat{a}^{\dagger}|N-1\rangle=0$. $\hat{a}^{\dagger}\hat{a}=\sum_{m,n}n\delta_{m,n}|m\rangle\langle n|=\sum_{n=0}^{N-1}n|n\rangle\langle n|=\text{diag}(0,1,2,\dots,N-1)$, $\hat{a}\hat{a}^{\dagger}=\sum_{n=0}^{N-2}(n+1)|n\rangle\langle n|=\text{diag}(1,2,\dots,N-1,0)$. The commutator becomes $\hat{a},\hat{a}^{\dagger}]=\text{diag}(1,1,\dots,1,-(N-1))$. The expectation value at $t=0$ is
$\mathbf{D}_{ii}(0)=\text{Tr}\{\hat{\rho}_{SS;i}[\hat{a}_{i},\hat{a}_{i}^{\dagger}]^{\dagger}[\hat{a}_{i},\hat{a}_{i}^{\dagger}]\}=\sum_{n=0}^{N-1}\langle n|\hat{\rho}_{SS;i}[\hat{a}_{i},\hat{a}_{i}^{\dagger}]^{\dagger}[\hat{a}_{i},\hat{a}_{i}^{\dagger}]|n\rangle=\sum_{n=0}^{N-2}P_{n}+P_{N-1}(N-1)^{2}=(1-P_{N-1})+(N-1)^{2}P_{N-1}=1+8P_{3},$ (45)
where $P_{n}=\langle n|\hat{\rho}_{SS;i}|n\rangle={\rm Tr}[\hat{\rho}_{SS;i}|n\rangle\langle n|]=\langle n|\text{Tr}_{j\neq i}[\hat{\rho}_{SS}]|n\rangle$, and thus $P_{N-1}$ is the population of the highest Fock state $|N-1\rangle=|3\rangle$ at site i. The identity OTOC can be recovered by increase $N_{max}$ until $P_{N_{max}}\to 0$.
In GMF solver, each site $l$ is solved using a local Liouvillian $\mathbf{L}_{l}$ where neighbors are treated as constant $c$-numbers ($\Psi_{l\pm 1}$) and the perturbation never spread along the chain. As shown in Fig.2(a), $D_{ii}(0)\neq 0$ and $\hat{a},\hat{a}^{\dagger}]=\text{diag}(1,1,1,-3)$. The squared commutator is $\text{diag}(1,1,1,9)$. The OTOC at site 3 decays to zero at long time, implying that the information of initial perturbation is lost to the environment quickly due to the dissipation. Sites 1 and 2 have much lower dissipation ($\gamma=0.1$) and the damped nonlinear oscillations represent the effects of coherent tunneling ($J$) and the nonlinear energy shifts ($U$). The OTOC tracks how information of initial $a^{\dagger}$ operator is scrambled across the local Hilbert space before dissipation eventually takes over. For stable (nonchaotic) NESS, the Heisenberg operator $\hat{a}(t)$ eventually loses its operator character and converges to a $c$-number steady-state value $\langle a\rangle_{ss}$, such that $\langle a\rangle_{ss}\mathbb{1},\hat{a}^{\dagger}]=0$. Any initial perturbation from $\hat{a}_{1}^{\dagger}$ is eventually dampened by the dissipation and the sensitivity $\partial a(t)/\partial a(0)\sim e^{-\gamma t/2}$ vanishes as $t\to\infty$. This is in contrast with the chaotic regime where the OTOC would saturate at a high value ($\sim 2n^{2}$) rather than returning to zero.
For exact solver as shown in Fig.2(b), using Eq.45, for $F=2.5$, $D_{1,1}(0)\approx 1.9$ implies for site 1 $P_{N-1}\approx(D_{1,1}(0)-1)/8=0.11$, reflecting the local quantum fluctuations. Here the deviation for $D_{1,1}(0)$ from 1 is due to the strong drive which push the occupation toward the cutoff value. Sites 2 and 3 start at 0 because $\hat{a}_{2,3}(0),\hat{a}_{1}^{\dagger}(0)]=0$ due to spatial commutation relations. The subsequent rise $D_{2,2},D_{3,3}\sim(Jt)^{2}$ reflects the hopping-induced ballistic or perturbative operator spreading due to the information scrambling[23, 24, 25]. The eventual decay to zero is the signature of a stable NESS, where local dissipation removes the phase memory of the initial perturbation. At short times the Hamiltonian (coherent) part of the dynamics dominates over the dissipators. Even in a stable system, the local operator $\hat{a}_{1}(0)$ begins to spread via tunneling $J$. As $\hat{a}_{1}(t)$ evolves, it becomes a multi-body operator with nonzero commutator $\hat{a}_{l}(t),\hat{a}_{j}(0)]$. The OTOC increases as the operator grows in complexity and spatial extent. This scrambling is a measure of how information becomes hidden in multi-body correlations before the environment has a chance to dissipate it.
The system enters into the nonlinear regime for lower dissipation ($J,U\gg\gamma$) and higher $F$ to pushes the system into highly excited states where the nonlinearity $U$ becomes more dominant.
In linear regime the decay of OTOC is dominated by dissipative damping (loss channel) and finite Liouvillian gap. is the global decoherence where particles and information are removed by the dissipators, and the perturbation propagates like a coherent wave through the chain and flows out of the drain site. For chaotic regime the the decay cause is the local decoherence (scrambling) where the information is hidden in high-order multi-site correlations. The nonlinearity $U$ causes trajectories to diverge exponentially. The perturbation is randomized across the entire Hilbert space. In the intermediate dephased nonlinear phase, there is dissipation-induced thermalization (local phase randomization due to the dissipation and interaction-induced dephasing) and the perturbation is suppressed by dissipative damping before it can evolve into many-body complexity. OTOC decays toward zero in a stable NESS and operator spreading is followed by strong interaction-induced dephasing and phase diffusion.
The exponential growth of OTOC in chaotic regime is related to the dynamical instability where the quantum memory is hidden by the complexity of many-body degrees of freedom instead of removed by the drain. Information exponentially populating the entire Hilbert space. However, in strongly dissipative Lindblad dynamics with a stable NESS, the gapped Liouvillian spectrum is contractive (negative real part) without positive Lyapunov exponent, where the OTOC exhibit a rapid short-time increase due to operator spreading, but does not display genuine exponential growth.
V.2 nonlocal OTOC
For nonlocal case, the commutator $\hat{a}_{i}(t),\hat{a}_{j}^{\dagger}(0)]$ measures how much a perturbation applied to site $j$ initially fails to commute with a measurement performed at site $i$ at $t$. In linear and stable regime, the growth of OTOC is bounded or oscillatory and the OTOC may exhibit a rapid short-time increase due to operator spreading. In chaotic regime, the norm of this commutator grows exponentially as $\sim e^{\lambda_{L}t}$, where $\lambda_{L}$ is the positive Lyapunov exponent. Thus chaos is defined as the exponential instability of trajectories within the NESS manifold. The exponential growth of OTOCs is commonly associated with dynamical instability and quantum chaos in closed or weakly dissipative systems. In chaotic regime of the open system, the OTOC saturates at $D(t\to\infty)\sim 2\langle n\rangle^{2}$. The memory of the initial perturbation is randomized (scrambled) across all degrees of freedom, but it isn’t lost to the environment.
For non-Hermitian operator $a_{i}(t),a^{{\dagger}}_{j}(0)]$, the operator OTOC can be written as
$\mathbf{D}_{ij}(t)=\text{Tr}\left[\hat{\rho}_{SS}[a_{i}(t),a^{{\dagger}}_{j}(0)]^{{\dagger}}[a_{i}(t),a^{{\dagger}}_{j}(0)]\right]=\langle[\hat{a}_{i}(t),\hat{a}_{j}^{\dagger}(0)]^{\dagger}[\hat{a}_{i}(t),\hat{a}_{j}^{\dagger}(0)]\rangle_{SS}$ (46)
$=\langle\hat{a}_{j}(0)\hat{a}_{i}^{{\dagger}}(t)\hat{a}_{i}(t)\hat{a}_{j}^{\dagger}(0)-\hat{a}_{j}(0)\hat{a}_{i}^{\dagger}(t)\hat{a}_{j}^{\dagger}(0)\hat{a}_{i}(t)-\hat{a}_{i}^{\dagger}(t)\hat{a}_{j}(0)\hat{a}_{i}(t)\hat{a}_{j}^{\dagger}(0)+\hat{a}_{i}^{\dagger}(t)\hat{a}_{j}(0)\hat{a}_{j}^{\dagger}(0)\hat{a}_{i}(t)\rangle$
$=2\langle\hat{a}_{i}^{\dagger}(t)\hat{a}_{i}(t)\rangle(\langle\hat{a}_{j}^{\dagger}(0)\hat{a}_{j}(0)\rangle+1)-2\text{Re}(\langle\hat{a}_{i}^{\dagger}(t)\hat{a}_{j}(0)\hat{a}_{i}(t)\hat{a}_{j}^{\dagger}(0)\rangle)$
where in GMF framework we have $\langle\hat{a}_{j}(0)\hat{a}_{i}^{{\dagger}}(t)\hat{a}_{i}(t)\hat{a}_{j}^{\dagger}(0)\rangle=\langle\hat{a}_{j}(0)\hat{a}_{j}^{\dagger}(0)\rangle\langle\hat{a}_{i}^{{\dagger}}(t)\hat{a}_{i}(t)\rangle=(\langle n_{j}\rangle+1)\langle n_{i}\rangle$, $\langle\hat{a}_{i}^{\dagger}(t)\hat{a}_{j}(0)\hat{a}_{j}^{\dagger}(0)\hat{a}_{i}(t)\rangle=\langle\hat{a}_{i}^{\dagger}(t)\hat{a}_{i}(t)\rangle\langle\hat{a}_{j}(0)\hat{a}_{j}^{\dagger}(0)\rangle=\langle n_{i}\rangle(\langle n_{j}\rangle+1)$. In chaotic regime, where the drive $F$ and $n$ are large, a high value of $\mathbf{D}_{ij}$ might simply reflect a high density rather than the quality of scrambling since $\mathbf{D}_{ij}$ scales with the photon number $\langle\hat{n}\rangle$. Thus the bracket in above expression accounts for the NESS with conserved $\langle\hat{n}\rangle$. When the system reaches its maximum scrambling state (long-time limit), all causal phase correlations between site $i$ and site $j$ are lost. Assuming $\langle\hat{a}_{i}\rangle\to 0$ in the chaotic bulk, the four terms in the expansion of $\mathbf{D}_{ij}$ simplify. Only the terms representing incoherent classical correlation survive, i.e., $\langle\hat{a}_{j}(0)\hat{a}_{i}^{\dagger}(t)\hat{a}_{i}(t)\hat{a}_{j}^{\dagger}(0)\rangle$ and $\langle\hat{a}_{i}^{\dagger}(t)\hat{a}_{j}(0)\hat{a}_{j}^{\dagger}(0)\hat{a}_{i}(t)\rangle$. The incoherent terms stabilize at NESS where the density at site $i$ is no longer fluctuating because the system has reached a balance between the drive and the dissipation. While the coherent cross-terms (representing the transport of phase information and nonlocal quantum correlation/interference between two sites) vanish, i.e., $\hat{a}_{j}(0)\hat{a}_{i}^{\dagger}(t)\hat{a}_{j}^{\dagger}(0)\hat{a}_{i}(t)-\hat{a}_{i}^{\dagger}(t)\hat{a}_{j}(0)\hat{a}_{i}(t)\hat{a}_{j}^{\dagger}(0)=0$, thus $\mathbf{D}_{ij}(t\to\infty)\approx 2\langle n_{i}\rangle_{SS}(\langle n_{j}\rangle_{SS}+1)$, which represents the total available fluctuation power between the two sites. This confirms that the scrambling has converted all coherent signal into incoherent noise intensity distributed across the sites. The phase information is scrambled across many degrees of freedom such that $\hat{a}_{i}(t),\hat{a}_{j}(0)]\neq 0$ with Heisenberg evolution $\hat{a}_{i}(t)=u(t)\hat{a}_{i}(0)+v(t)\hat{a}_{j}(0)$ and the expectation value vanishes, $\langle\hat{a}_{i}(t)\rangle=0$. This decay rate is fundamentally linked to the Lyapunov exponent $\lambda_{L}$ such that $|[\hat{a}_{i}(t),\hat{a}_{j}(0)]|^{2}\sim e^{\lambda_{L}t}$. Note that here the $\lambda_{L}$ decay is the signature of the internal chaotic scrambling where the information (quantum memory) is destroyed globally by the chaos, which a local observer experiences as local decoherence (instead of the global decoherence due to the drive and drain)[21].
The OTOC measures sensitivity of the field at site $i$ to a perturbation at site $j$ at $t=0$, probing the scrambling of phase information across the chain. The normalized OTOC (phase OTOC) reads
$\tilde{\mathbf{D}}_{ij}(t)=\frac{2\langle\hat{a}_{i}^{\dagger}(t)\hat{a}_{i}(t)\rangle(\langle\hat{a}_{j}^{\dagger}(0)\hat{a}_{j}(0)\rangle+1)-2\text{Re}(\langle\hat{a}_{i}^{\dagger}(t)\hat{a}_{j}(0)\hat{a}_{i}(t)\hat{a}_{j}^{\dagger}(0)\rangle)}{2\langle\hat{a}_{i}^{\dagger}(t)\hat{a}_{i}(t)\rangle\langle\hat{a}_{j}(0)\hat{a}_{j}^{{\dagger}}(0)\rangle}$ (47)
$=1-\text{Re}\left[\frac{\langle a_{i}^{\dagger}(t)a_{j}a_{i}(t)a^{{\dagger}}_{j}\rangle_{SS}}{\langle a_{i}^{\dagger}a_{i}\rangle_{SS}\langle a_{j}a^{{\dagger}}_{j}\rangle_{SS}}\right].$
In this normalized form, $\tilde{\mathbf{D}}\to 0$ means the phases are perfectly correlated (regular), $\tilde{\mathbf{D}}\to 1$ means the phase information is completely scrambled (chaotic/Thermal). Cauchy-Schwarz inequality $|\text{Tr}(\hat{\rho}\hat{A})|\leq||\hat{\rho}||_{2}||\cdot||_{2}$ (where $||\cdot||_{2}$ is the Hilbert-Schmidt norm) implies, since $\hat{\rho}||_{2}=\sqrt{\text{Tr}(\rho^{2})}\leq 1$ (with equality only for pure states), the OTOC is bounded by the norm of the squared commutator[20]. In a finite Hilbert space the operator OTOC cannot grow indefinitely. The Cauchy-Schwarz inequality, provides the theoretical maximum growth. Saturation occurs when the operator $\hat{a}_{i}(t)$ becomes uniformly distributed over the available operator basis.
[…]
In the weak drive/linear limit, commutators decay or remain constant. $D_{1,L}(\tau\to\infty)\approx 0$. While in the strong coupling regime ($U\gg\gamma$), the non-linearity scrambles local information globally.
@@
The OTOC exhibits characteristic exponential growth at early times, $D(\tau)\sim e^{\lambda_{L}\tau}$ (where $\lambda_{L}$ is the quantum Lyapunov exponent), before saturating to a finite value $O(1)$ determined by the local Hilbert space dimension. This saturation indicates that information about the drive at site 1 has been effectively "forgotten" or scrambled across the entire many-body Hilbert space of the chain.
@@
Fig.2 show the OTOC as a function of time in unit of inverse hopping amplitude ($1/J$; such that the value $t=1$ corresponds to the characteristic time required for a particle to tunnel between neighboring sites). for site 3 (drain site) is of the strong dissipative and quasilinear regime, Although quantum information propagates to this site as indicated by the initial rise, it is quickly lost to the environment due to strong coupling with the bath. The vanishing saturation value is consistent with a strong dissipation or quasilinear regime, where information is not scrambled locally but rather "flushed out" before it can accumulate.